# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v1

@@ VII-B2 Model introspection
Literature that evaluates models by correlating model uncertainty with error is, comparatively, much more thorough and standardized. This is especially evident by the frequent use of the PAvPU metric. Furthermore, a majority of the available research use Monte Carlo Dropout and correlated this to the output entropy to the prediction accuracy. The disadvantage of relying on MC Dropout when the uncertainty is crucial, has been discussed in Section VII-A2. Therefore, we recommend incorporating explicit Variational Inference at specific points in the network to achieve more accurate uncertainty quantification. These points could be selected based on the layers that most significantly influence the output segmentation.
@@ VII-B3 Model generalization
Epistemic uncertainty quantification methods can result in improved performance as it often serves as a surrogate for model combination. Therefore, uncertainty quantification is can readily be available in the toolbox of deep learning researchers together with other commonly used regularization techniques. In this case, simple methods such as ensembling and MC dropout are despite their criticisms relatively harmless. Furthermore, parallels can be made with MC Dropout and placing an L2-norm penalty on the model weights. While, this benefits is compelling, improved model performance can also be obtained with more computationally efficient regularizers. Therefore, improved performance should not be the end goal of uncertainty quantification, but considered as an ancillary advantage.
@@ VII-B4 Active Learning
Active Learning, while being a challenging task, holds significant promise by potentially reducing the need for intensive labeling procedures that often require specialized expertise. Furthermore, this kind of approach hints towards a strong collaboration between humans (often referred to as an external oracle) and Artificial Intelligence, which can accelerate adoption of such models in sectors requiring extensive specialization. Furthermore, Active Learning can accelerate privacy-centric collaboration when combined in a federating setting, enabling the active improvement of safety-critical models with human-in-the-loop intervention across local modals.
However, active learning requires a well-generalized model on little data. This is a challenging task in the realm of deep learning-based segmentation, which usually deals with high-dimensional data and often requires datasets of substantial size. Furthermore, many uncertainty-based approaches often extend traditional active learning for classification by simply aggregating the pixel’s metrics together. For imbalanced settings, it has been shown that it can perform even worse than random selection [215]. Therefore, additional modifications such as target, boundary or diversity awareness [198, 203], or region-based annotating [202, 216] are often required to apply Active Learning to segmentation tasks.
@@ VII-C Future work
In this penultimate section, we will provide point-by-point recommendations for future work.
@@ VII-C1 Exploring generative models
Generative modeling, a rapidly evolving field, has successfully been employed for quantifying observer variability. The benefits of quickly applying its developments to segmentation problems have been evident with state-of-the-art DDPMs [90], which were initially proposed for unsupervised image generation. Also, literature on unsupervised VAEs have greatly benefited the PU-Net [65, 79, 70]. Therefore, we advocate for a deeper contextualization of contemporary research within probabilistic segmentation models. In fact, any unsupervised generative model can theoretically be translated to the supervised setting through intricate conditioning and can therefore be used as a probabilistic segmentation model.
Given this flexibility, it remains unclear why Normalizing Flow (NF) remains underutilized for this task. Specifically continuous NFs, which approximate the time-dependent score function, strongly resembles DDPMs which have extensively been used for segmentation problems. Furthermore, NFs enable explicit and exact evaluation of the likelihood, which can aid in further interpretation of model predictions. Instead of modeling an iterative stochastic linear Gaussian process, NFs construct a series of invertible functions towards an isotropic Gaussian and inference times are therefore much faster than that of DDPMs. It should be noted, however, that the invertibility constraint greatly hinders expressivity of intermediate functions, leading to memory-intensive architectures.
@@ VII-C2 Variational inference and beyond
A valuable contribution to the field pertains a comprehensive benchmark paper, which compares all available epistemic uncertainty quantification methods across a wide range of datasets. In particular, such study can elucidate the data-dependent preference for specific methodologies (i.e. why ensembling or MC Dropout is often preferred over explicit VI). Additionally, recent studies have shown the benefits of moving from a few large to many small experts for when using a MoE ensemble (see Section V-C for language modeling [217]. Future work should also experiment with this.
Approaches besides VI, such as Markov Chain Monte Carlo (MCMC) or Laplace Approximations, are also viable options to approximate the Bayesian posterior. Especially the Laplace approximation can be very beneficial, as it is easily applicable to pretrained networks. Notably, both Laplace approximation and VI are biased and operate in the neighborhood of a single mode, while MCMC methods are useful when expecting to fit multi-modal parameter distributions. To the best of our knowledge, these approaches have not been studied within the context of uncertainty quantification in segmentation.
@@ VII-C3 Single-pass uncertainty
The multiple forward passes required in Bayesian uncertainty quantification can incur cumbersome additional costs. Hence, considerable efforts have been made towards deterministic uncertainty models [218, 219, 220], which only depend on a single forward pass. Mukhoti et al. [218] show that Gaussian Discriminant Analysis after training with SoftMax predictive distribution can in some instances surpass methods such as MC Dropout and ensembling. This approach achieves faster computation, while also providing both epistemic and aleatoric uncertainty.
Along similar lines, Evidential Deep Learning also possess the advantage of quantifying both uncertainties with a single forward pass. This framework is based on a generalization of Bayes theorem, known as the Dempster-Shafer Theory of Evidence (DST) [221]. While common in Bayesian probability, DST does not require prior probabilities and bases subjective probabilities on belief masses assigned on a frame of discernment, i.e. the set of all possible outcomes. The use of Evidential Deep Learning has seen success in conventional classification problems [222], and Ancha et al. [223] recently applied this concept to segmentation to decouple aleatoric and epistemic uncertainty within a single model. Unfortunately, there has not been much research beyond this.
@@ VII-C4 Improved CNN architectures
With the introduction of vision transformers [224], CNN-based models have been challenged in the task of semantic segmentation [225, 226, 227]. Regardless the success of vision transformers across many domains, it is clear that CNN-based encoder-decoder models such as the U-Net remain the preferred backbone [39]. This is mainly because CNNs already possess desirable inductive biases, while transformers conversely require extensive pretraining with large datasets [228]. Nonetheless, CNNs have benefited from the recent developments in transformers. For instance, “ConvNext” takes inspiration from contemporary transformers to modernize existing ResNet-based CNNs, retaining the inductive biases of convolutional filters and achieving significant performance gains [229]. Since many innovations in this field focus on the technique of uncertainty quantification, used backbones receive less attention and might often be outdated. Our recommendation is to improve current models with developments in general CNN-based architectures.
@@ VII-C5 Distributing-free modeling
A distribution-free framework known as Conformal Deep Learning produces prediction sets that guarantee to contain the ground-truth with a user-defined probability. With the help of an additional calibration set, a heuristic notion of ambiguity (i.e. miscalibrated softmax outputs) is transformed to rigorous uncertainty and is especially renowned for being model-agnostic, simple and highly flexible [230]. Very recently, conformal prediction has been applied to segmentation problems [231, 232, 233, 234], thereby enjoying the aforementioned benefits of this framework and indicating an increased traction. We recommend further research in this direction to discover novel applications and to benchmark against current architectures.
@@ VII-C6 Volumetric segmentation
In many clinical datasets, the volumetric data is in most instances sliced to patches and processed with conventional 2D-based CNN-models. However, such models are often straightforward extensions of existing 2D models (and almost exclusively VAE-based), rarely addressing novel challenges introduced with additional dimensionalities. For example, 3D extensions of the PU-Net simply remain to use similar techniques to insert latent samples in the decoding networks. Also, we have noted that works related to 3D BNN training require group normalization and KL-annealing for accurate generalization. Therefore, general guidelines related to translating 2D segmentation models to 3D can be of great benefit for practitioners looking to implement models for volumetric segmentations problems.
Also, methodologies need to be developed in order to appropriately compare 2D to 3D models. For example, the GED can not be estimated volumetrically with 2D models. Furthermore, translating already computationally intensive models (e.g. based on DDPMs or PixelCNNs) to three dimensions is another huge challenge due to the memory requirement involved volumetric data, which will even increase the training load and inference time even further. Fortunately, research dedicated to volumetric segmentation is prevalent and its successes underlines the need of more investigation related to three-dimensional uncertainty quantification, which more closely aligns with real-world clinical practice and therefore encourages faster adoption.
@@ VIII Conclusion
Modeling the uncertainty of segmentation models is essential for accurately assessing the reliability in their predictions. With the vast body of literature, encompassing diverse applications and modalities, the need for a comprehensive and systematic overview of the field is addressed. We present clear definitions and notation of methodologies that attempt uncertainty modeling by considering the field from a theoretical perspective and relating this to various pertinent applications. Aleatoric uncertainty can be modeled in pixel or latent space with generative models, or implicitly be expressed with test-time augmentation. Epistemic uncertainty is captured with Variational Inference on the parameter distribution, or approximated with Monte Carlo Dropout or model ensembling. Our findings show that both aleatoric and epistemic uncertainty modeling enable the practice of four distinct downstream tasks. Which in turn enable us to highlight the main challenges and limitations of current work, both related to the theoretical frameworks and real-world applications.
Our future recommendation pertain aligning the field with advancements in general generative modeling and deep learning architectures. Furthermore, the adoption of deterministic uncertainty quantification methods that do no require multiple forward passes such as Conformal and Evidential Deep Learning is suggested. Especially the latter approach is interesting due to its ability to encapsulate and express both uncertainty types. Since most epistemic uncertainty quantification is performed with approximate Variational Inference, a comprehensive benchmark study of these techniques as well as exploration of other techniques such as Markov Chain Monte Carlo (MCMC) or Laplace Approximation is a beneficial contribution to the field. Finally, due to the clinical relevancy of uncertainty in semantic segmentation, more attention to models catered to volumetric data is advised. In this manner, the review paper guides researchers on the topic of probabilistic segmentation and suggests future endeavors within the lightning-fast evolving field of Deep Learning-based Computer Vision.

### v3

@@ VII-B2 Model introspection
Literature that evaluates models by correlating model uncertainty with error is, comparatively, much more thorough and standardized. This is especially evident by the frequent use of the PAvPU metric. Furthermore, a majority of the available research use MC Dropout and correlate the output entropy to the prediction accuracy. The disadvantage of relying on MC Dropout when the uncertainty is crucial, has been discussed in Section VII-A2. Therefore, we recommend incorporating explicit VI at specific points in the network to achieve more accurate uncertainty quantification. These points can be selected based on the layers that influence the output segmentation most significantly.
@@ VII-B3 Model generalization
Epistemic uncertainty quantification methods can result in improved performance, as it often serves as a surrogate for model combination. Therefore, uncertainty quantification can be readily available in the toolbox of deep learning researchers together with other commonly used regularization techniques. In this case, simple methods such as ensembling and MC dropout, are despite their criticisms relatively harmless. Furthermore, a parallel characterization can be made with MC Dropout and placing an $L_{2}$-norm penalty on the model weights. While this benefit is compelling, improved model performance can be also obtained with more computationally efficient regularizers. Therefore, improved performance should not be the ultimate goal of uncertainty quantification, but considered as an ancillary advantage.
@@ VII-B4 Active Learning
The challenging task of Active Learning holds a significant promise and can potentially reduce the need for intensive labeling procedures, which often require specialized expertise. Furthermore, this kind of approach points towards a strong collaboration between humans (often referred to as an external oracle) and Artificial Intelligence, which can accelerate the adoption of such models in sectors requiring extensive specialization. Furthermore, Active Learning can accelerate privacy-centric collaboration when combined with a federated setting, enabling the active improvement of safety-critical models with human-in-the-loop intervention across local models.
However, Active Learning requires a well-generalized model using little data. This is a challenging task in the realm of deep learning-based segmentation, which usually deals with high-dimensional data and typically requires datasets of substantial size. Furthermore, many uncertainty-based approaches regularly extend conventional Active Learning for classification to segmentation by simply aggregating the pixel metrics together. For imbalanced settings, it has been shown that it can perform even worse than random selection [217]. Therefore, additional modifications such as target, boundary or diversity awareness [200, 205], or region-based annotation [204, 218] are often required to apply Active Learning to segmentation tasks.
@@ VII-C Future work
In this penultimate section, we will provide recommendations for future work.
@@ VII-C1 Exploring generative models
Generative modeling, a rapidly evolving field, has successfully been employed for quantifying observer variability. The benefits of quickly applying its developments to segmentation problems have been evident with state-of-the-art DDPMs [90], which were initially proposed for unsupervised image generation. Also, literature on unsupervised VAEs have greatly benefited the PU-Net [73, 63, 68]. Therefore, we advocate for a deeper contextualization of contemporary generative modeling research within probabilistic segmentation models. In fact, any unsupervised generative model can be theoretically translated to the supervised setting through intricate conditioning and can be therefore used as a probabilistic segmentation model.
Given this flexibility, it remains unclear why Normalizing Flows (NFs) remain underutilized for this task. Specifically continuous NFs, which approximate the time-dependent score function, strongly resemble DDPMs which have been extensively used for segmentation problems. Furthermore, NFs enable explicit and exact evaluation of the likelihood, which can aid in further interpretation of model predictions. Instead of modeling an iterative stochastic linear Gaussian process, NFs construct a series of invertible functions towards an isotropic Gaussian, so that inference times are therefore much faster. However, it should be noted that the invertibility constraint greatly hinders the expressiveness of intermediate functions, leading to memory-intensive architectures.
@@ VII-C2 Variational Inference and beyond
A valuable contribution to the field pertains a comprehensive benchmark paper, which compares all available epistemic uncertainty quantification methods across a wide range of datasets. In particular, such study can elucidate the data-dependent preference for specific methodologies (i.e. why ensembling or MC Dropout is often preferred over explicit VI). Additionally, recent studies have shown the benefits of moving from a few large, to many small experts when using an MoE ensemble (see Section V-C) for language modeling [219]. Future work should also experiment with this.
Approaches besides VI, such as Markov Chain Monte Carlo (MCMC) or Laplace Approximations, are also viable options to approximate the Bayesian posterior. Especially the Laplace Approximation can be very beneficial, as it is easily applicable to pretrained networks. Notably, both Laplace Approximation and VI are biased and operate in the neighborhood of a single mode, while MCMC methods are a more attractive option when expecting to fit multi-modal parameter distributions. To the best of our knowledge, these approaches have not been studied within the context of uncertainty quantification in segmentation.
@@ VII-C3 Single-pass uncertainty
The multiple forward passes required in many Bayesian uncertainty quantification can incur cumbersome additional costs. Hence, considerable efforts have been made towards deterministic uncertainty models [220, 221, 222], which only depend on a single forward pass. Mukhoti et al. [220] show that Gaussian Discriminant Analysis after training with SoftMax predictive distribution can in some instances surpass methods such as MC Dropout and ensembling. This approach achieves faster computation, while also providing both epistemic and aleatoric uncertainty.
Along similar lines, Evidential Deep Learning also possesses the advantage of quantifying both uncertainties with a single forward pass. This framework is based on a generalization of Bayes theorem, known as the Dempster-Shafer Theory of Evidence (DST) [223]. While common in Bayesian probability, DST does not require prior probabilities and bases subjective probabilities on belief masses assigned on a frame of discernment, i.e. the set of all possible outcomes. The use of Evidential Deep Learning has seen success in conventional classification problems [224], and Ancha et al. [225] recently applied this concept to segmentation to decouple aleatoric and epistemic uncertainty within a single model. Unfortunately, further research has not been published beyond this.
@@ VII-C4 Improved CNN architectures
With the introduction of Vision Transformers [226], CNN-based models have been challenged in the task of semantic segmentation [227, 228, 229]. Regardless the success of Vision Transformers across many domains, it is clear that CNN-based encoder-decoder models such as the U-Net, remain the preferred backbone architecture [39]. This is mainly because CNNs already possess desirable inductive biases, while Transformers conversely require extensive pretraining with large datasets [230]. Nonetheless, CNNs have benefited from the recent developments in Transformers. For instance, “ConvNext” takes inspiration from contemporary Transformers to modernize existing ResNet-based CNNs, retaining the inductive biases of convolutional filters and achieving significant performance gains [231]. Since many innovations in this field focus on the technique of uncertainty quantification, the applied backbones receive less attention and might gradually become outdated. Our recommendation suggests improving current models with developments in generic CNN-based architectures.
@@ VII-C5 Distributing-free modeling
A distribution-free framework known as Conformal Deep Learning, produces prediction sets that guarantee to contain the ground truth with a user-defined probability. With the help of an additional calibration set, a heuristic notion of ambiguity (i.e. miscalibrated softmax outputs) is transformed to rigorous uncertainty. This technique is especially renown for being model-agnostic, simple and highly flexible [232]. Quite recently, Conformal Prediction has been applied to segmentation problems [233, 234, 235, 236], thereby enjoying the aforementioned benefits and showing an increased traction of this framework. We recommend further research in this direction to discover novel applications and benchmark against current architectures.
@@ VII-C6 Volumetric segmentation
In many clinical datasets, the volumetric data are in most instances sliced to patches and processed with conventional 2D-based CNN-models. However, such models are often straightforward extensions of existing 2D models (and almost exclusively VAE-based), rarely addressing novel challenges introduced with additional dimensionalities. For example, 3D extensions of the PU-Net simply remain to use similar techniques for inserting latent samples in the decoding networks. Also, we have noted that works related to 3D BNN training require group normalization and KL-annealing for accurate generalization [179]. Therefore, general guidelines related to translating 2D segmentation models to 3D can be of great benefit for practitioners considering to implement models for volumetric segmentations problems.
Furthermore, methodologies need to be developed in order to appropriately compare 2D to 3D models. For example, the GED cannot be estimated volumetrically with 2D models. Moreover, translating the already computation-intensive models (e.g. based on DDPMs or PixelCNNs) to 3D is another challenge, due to the memory requirement involved with volumetric data, which will increase the training load and inference time even further. Fortunately, research dedicated to volumetric segmentation is prevalent and its successes underline the need of more investigation related to 3D uncertainty quantification, that more closely aligns with real-world clinical practice and therefore encourages faster adoption.
@@ VIII Conclusion
Modeling the uncertainty of segmentation models is essential for accurately assessing the reliability in their predictions. With the vast body of literature, encompassing diverse applications and modalities, the need for a comprehensive and systematic overview of the field is addressed. We present clear definitions and unified notation of methodologies that attempt uncertainty modeling, by considering the field from a theoretical perspective. Aleatoric uncertainty can be modeled in pixel or latent space with generative models, or implicitly be expressed with test-time augmentation. Epistemic uncertainty is captured with Variational Inference on the parameter distribution, or approximated with Monte Carlo Dropout or model ensembling. Our findings show that both aleatoric and epistemic uncertainty modeling enable the practice of four distinct downstream tasks, which enable to highlight the main challenges and limitations of current work, both related to the theoretical frameworks and to the real-world applications.
Our future recommendations pertain aligning the field with advancements in general generative modeling and Deep Learning architectures. Furthermore, the adoption of deterministic uncertainty quantification methods that avoid multiple forward passes, such as Conformal and Evidential Deep Learning, are suggested. Particularly the latter approach is interesting due to its ability to incorporate both uncertainty types. Since most epistemic uncertainty quantification is performed with approximate Variational Inference, a comprehensive benchmark study of these techniques, as well as exploration of other techniques such as Markov Chain Monte Carlo (MCMC) or Laplace Approximation, is a beneficial contribution to the field. Finally, due to the clinical relevance of uncertainty in semantic segmentation, more attention is advised to models catering to volumetric data. In this manner, the review paper guides researchers on the topic of probabilistic segmentation and suggests future endeavors within the lightning-fast evolving field of Deep Learning-based Computer Vision.